Model Summary Adjusted R Square Std. Error of the Estimate Model R Square 205 a. Predictors: (Constant), HIGHEST YEAR SCHOOLInvestigate the relationship between the respondent’s education (EDUC) and the education received by his or her father and mother (PAEDUC and MAEDUC, respectively). Calculate the correlation coefficient, the coefficient of determination, and the regression equation predicting the respondent’s education with father’s education only. Interpret your results. Determine the multiple correlation coefficient, the multiple coefficient of determination, and the regression equation predicting the respondent’s education with father’s and mother’s education. Interpret your results. Did taking into account the respondent’s mother’s education improve our prediction? Discuss this on the basis of the results from 5b. Using the regression equation from 5a, calculate the predicted number of years of education for a person with a father with 12 years of education. Then, repeat this procedure, adding in a mother’s 12 years of education and using the regression equation from 5b. Review the ANOVA results. Can you reject the null hypothesis that r2 = 0?
Answer #1

I can only answer the first part of the questions as the results of the regression with the mother's education are not given here.

The Correlation Coefficient = 0.452

Coefficient of Determination = r square = 0.452 * 0.452 = 0.2043

Regression Equation is

Y = 10.333+ 0.327 X

where Y = Dependent Variable = son' s education

X = Father's Education

Putting X =12 we get "

Y = 10.333 + 0.327 *12 = 14.257 years